OpenStudy (anonymous):

write an equation in standard form for the ellipse with center (0,0) vertex(5,0) covertex(0,-2)

OpenStudy (anonymous):

\[\huge \color{blue}{\frac{x^2}{25} +\frac{y^2}{4} = 1}\]